# Boter: Bootstrapping Knowledge Selection and Question Answering for Knowledge-based VQA

### v1

@@ 4.4. Qualitative Analysis
5. CONCLUSION
In this paper, we propose Boter, a novel framework to address the task of knowledge-based VQA. We adopt the multimodal language model to construct two modules: a Selector and an Answerer. The Selector selects key knowledge from the retrieved knowledge documents and the Answerer reasons on the key knowledge to answer questions. Through a cycle training method, we ensure consistent improvement of both modules’ abilities in knowledge selection and question answering. Compared with state-of-the-art methods, Boter achieves better performance on a challenging open-domain knowledge-based VQA benchmark (OK-VQA) and we conduct a comprehensive analysis to evaluate the effectiveness of our method.

### v3

@@ 4.4 Qualitative Analysis
5 Conclusion
In this paper, we propose a novel framework that leverages the large visual-language model to construct two modules: (1) Selector for finding key retrieved knowledge and (2) Answerer for reasoning on the knowledge to predict answers. We design a self-bootstrap learning method to improve their abilities, where the Selector chooses key knowledge documents for the Answerer and the Answerer provides pseudo-labels for the Selector. Compared with state-of-the-art methods, our method achieves better performance on a challenging open-domain knowledge-based VQA benchmark (OK-VQA) and we conduct a comprehensive analysis to evaluate the effectiveness of our method.
6 Acknowledgements
This work was supported by the Artificial Intelligence-National Science and Technology Major Project (2023ZD0121200), the National Natural Science Foundation of China (U21B2043, 62102416), and the Key Research and Development Program of Jiangsu Province under Grant BE2023016-3.
7 Limitations
Although our framework can effectively select key knowledge documents for answering question, it is inevitable that the knowledge still contains noise. In some cases, the model itself can answer the question without external knowledge, introducing extra knowledge may affect the performance. In the future, we can explore to dynamically select required knowledge to help itself answer questions.
In addition, there is a concern on the generalizability of the proposed method on other domains, especially when the initial DPR model can not retrieve gold standard context. In the future, we consider adopting a stronger multimodal retriever model to obtain more useful candidate knowledge documents, which enhances the generalizability of our framework.
Appendix A Appendix
Table 9: Performance comparison with state-of-the-art (SOTA) methods on the FVQA dataset.
Method Acc-1
Human 77.99
UnifER Guo et al. (2022) 55.04
FVQA Wang et al. (2017) 56.91
ZS-VQA Chen et al. (2021) 58.27
FVQA(Ensemble) Wang et al. (2017) 58.76
MM-Reasoner(Ensemble) Khademi et al. (2023) 61.10
Ours 63.3
Table 10: Performance comparison with state-of-the-art (SOTA) methods on the A-OKVQA dataset.
Method Direct Answer
val test
ClipCap Schwenk et al. (2022) 18.1 15.8
Pythia Jiang et al. (2018) 25.2 21.9
ViLBERT Lu et al. (2019) 30.6 25.9
LXMERT Tan and Bansal (2019) 30.7 25.9
KRISP Marino et al. (2021) 33.7 27.1
GPV-2 Kamath et al. (2022) 48.6 40.7
BLIP-2 T5-XL Li et al. (2023) 53.2 49.7
PromptCap + GPT-3 Hu et al. (2022b) 56.3 59.6
Ours 57.2 56.4
A.1 Experiments on Other Datasets.
We also evaluate our method on FVQA Fang et al. (2023) and A-OKVQA Schwenk et al. (2022) to demonstrate the effectiveness of our method. FVQA is a VQA dataset that mostly contains questions requiring external knowledge to answer, and provides supporting fact triplets alongside the image-question-answer triplets. A-OKVQA is an augmented successor of OK-VQA, containing 25K image-question pairs that require broader commonsense and world knowledge to answer. Due to A-OKVQA does not provide the knowledge source, we use Wikipedia Vrandečić and Krötzsch (2014) as the knowledge base.
As shown in Tab. 9, our method surpasses previous state-of-the-art methods, which demonstrates the effectiveness and generalization of our method. Tab. 10 shows the comparative results on the challenging A-OKVQA dataset. Our method achieved competitive results, which demonstrates the effectiveness of our method.
Table 11: Computational cost of our framework.
$K_{candidate}$ Memory (GB) Running Time (sec./sample)
10 21.3 1.0
15 21.4 1.1
30 22.7 1.3
A.2 Evaluation of Computational Cost
In Tab. 11, we show the computational cost of our framework using different knowledge candidates. As the number of candidate knowledge, the computational cost of our framework has a small increase.